Patti needs 8000 litres of a 35% solution of hydrogen chloride. How many litres of a 40% solution and how many litres of a 20% solution will she use?
1. How many litres of 40% solution & how many solutions of 20% solution?
2. Litre of 40% = x
Litre of 20% = y
x + y = 8000
.40x + .20y = .35
Many thanks for any assistance your able to share.
pmoss
From this point on I'm stuck..I don't know how to solve the problem.

Answer by rajagopalan(174)

Note * is used as multiplication symbol as x is used as the variable:
..
Let Litres of 40% = x
Then Litres of 20% = (8000-x)
But Total H Cl in 8000 litres of 35% solution = 0.35*8000= 2800 litres
H Cl in 40% solution =0.4*x
H Cl in 20% solution =0.2*(8000-x)
Now our equation is (0.4*x)+(0.2*(8000-x))=2800
0.4x+(8000-x)0.2=2800
0.4x-1600-0.2x=2800
0.2x-1600=2800
0.2x=2800-1600
0.2x=1200
x=1200/0.2=6000
so 40% solution = 6000 litres
and 20% solution =2000 litres
Okay